Household wiring often uses 2.0 mm diameter copper wires. The wires can get rather long as they snake through the walls from the fuse box to the farthest corners of your house.


What is the potential difference across a 16 m long, 2.0 mm diameter copper wire carrying a 7.3 A current?
Answer #1
Concepts and reason

The concepts required to solve this problem are resistivity and Ohm’s law.

First use the resistivity equation to solve for the resistance of the wire.

Later, calculate the potential difference by using the Ohm’s law.

Fundamentals

The resistivity of a wire is given as,

ρ=ALR\rho = \frac{A}{L}R

Here, LL is the length of the wire, AA is the area of cross section of the wire, and RR is the resistance of wire.

The Ohm’s law states that the potential difference in a resistor is directly proportional to the current flowing through the circuit keeping all other conditions constant. The constant of proportionality is called resistance. The equation of Ohm’s law is,

V=IRV = IR

Here, VV is the applied potential difference, II is the current, and RR is the resistance.

(A)

Use the resistivity equation to solve for the resistance.

Substitute πd24\pi \frac{{{d^2}}}{4} for AA in the equation ρ=ALR\rho = \frac{A}{L}R and solve for RR.

ρ=π(d24)LRR=4Lρπd2\begin{array}{c}\\\rho = \frac{{\pi \left( {\frac{{{d^2}}}{4}} \right)}}{L}R\\\\R = \frac{{4L\rho }}{{\pi {d^2}}}\\\end{array}

Use the Ohms law to solve for potential difference.

Substitute 4Lρπd2\frac{{4L\rho }}{{\pi {d^2}}} for RR in the equationV=IRV = IR.

V=4LρIπd2V = \frac{{4L\rho I}}{{\pi {d^2}}}

Substitute 1.68×108Ωm1.68 \times {10^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}} forρ\rho , 7.3A{\rm{7}}{\rm{.3 A}} forII, 16m16{\rm{ m}} forLL, and 2.0mm{\rm{2}}{\rm{.0 mm}} for dd in the equation V=4LρIπd2V = \frac{{4L\rho I}}{{\pi {d^2}}} and calculate the potential difference.

V=4(1.68×108Ωm)(7.3A)(16m)π(2.0mm)2=4(1.68×108Ωm)(7.3A)(16m)π(2.0mm(103m1mm))2=0.625V\begin{array}{c}\\V = \frac{{4\left( {1.68 \times {{10}^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}}} \right)\left( {{\rm{7}}{\rm{.3 A}}} \right)\left( {16{\rm{ m}}} \right)}}{{\pi {{\left( {{\rm{2}}{\rm{.0 mm}}} \right)}^2}}}\\\\ = \frac{{4\left( {1.68 \times {{10}^{ - 8}}{\rm{ }}\Omega \cdot {\rm{m}}} \right)\left( {{\rm{7}}{\rm{.3 A}}} \right)\left( {16{\rm{ m}}} \right)}}{{\pi {{\left( {{\rm{2}}{\rm{.0 mm}}\left( {\frac{{{{10}^{ - 3}}{\rm{ m}}}}{{1{\rm{ mm}}}}} \right)} \right)}^2}}}\\\\ = 0.625{\rm{ V}}\\\end{array}

Ans:

The potential difference across the wire is0.625V{\rm{0}}{\rm{.625 V}}.